OpenStudy (anonymous):

a+1 over 6 = 2/a what is a

OpenStudy (anonymous):

\[(a+1)/6=2/a\] multiply both sides by 6a what would you get?

OpenStudy (anonymous):

6a^2+1 = 12a^2

OpenStudy (anonymous):

a(a+1)=12 a^2 +a-12=0 use quadratic formula to find a what did you get for a?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

or you can re-write it as: (a+4)(a-3)=0 (i used quadratic formula to find roots... honestly...) so, the answer would be: a=3 and a=-4

OpenStudy (anonymous):

are you good?